The velocity of the bullet reduces by (1/20)th of its initial velocity after passing through a plank. The maximum number of planks it can cross is
1. 11
2. 10
3. 19
4. 20

Grade:11

1 Answers

Arun
5 years ago

If the bullet looses 1/20th of its velocity then

V= 19/20u

Since, v² = u² + 2as

(19/20u)² = u² - 2as

2as = u² - (19/20u)² ................(1)

Now, initial velocity = 0

Therefore 0² = u² - 2a (ns)

u² = 2as (n)

u² = n [ u² - 19²/20² u²]

u² = nu² [ 20²-19²/20²]

20² = n [ 20² - 19² ]

20² = n [ (20+19)(20-19) ]

n = 400/39

n ~ 11 planks